Reports 1-1 of 1 Clear search Modify search
takahiro.yamamoto - 15:10 Monday 27 September 2021 (18354) Print this report
Control preparation for ITMY TWR


- Filterbanks were cleaned up for guardian.
- They were based on ones on k1visetmxt
- Set-points and matrices have been touched yet.


At first, some FMs which were blank were set as OFF in safe.snap (See Fig.1).

Integrators on BLEND for converting velocity and displacement were moved to ACCBLEND same as klog#18152 (See Fig.2).
Integrators haven't been turned on yet. It should be turned ON in proper timing for data acquisition (maybe maintenance day?).

ramp times were set as 20s in safe.snap in order to unify other TypeA suspensions (See Fig.3 and Fig.4)
There is no difference yet because offset values are set as 0.

FM positions of DC and DAMP were swapped in order to avoid glitches as reported in klog#18249.
And also, null filter was implemented on FM1 for floating.
Now null, DAMP and DC are set on FM1, FM3 and FM4, respectively (See Fig.5).

Because a new guardian turns ON/OFF opticalign by changing an OFFSET switch not a GAIN value,
GAIN and OFFSET were set as 1.0 and OFF in safe.snap (See Fig.6 and Fig.7).

A DAMP filter was moved from FM4 to FM3.
null and DC filter was newly prepared on FM1, and FM4, respectively.
Current implementation is shown in Fig.8 and Fig.9.

In order to avoid numerical rounding error reported in klog#18254,
DAMP filter was modified as shown in Fig.10 (0Hz zero -> 0.01Hz zero).
Current DAMP, DC, and DAMP+DC filters are represents as red curves. Blue one shows an original DAMP filter.
Images attached to this report
Search Help